OpenStudy (anonymous):

MEDALSSS!!! + FAN QUICK QUESTION A student scored 83 and 91 on her first two quizzes. Write and solve a compound inequality to find the possible values for a third quiz score that would give her an average between 85 and 90, inclusive.

sammixboo (sammixboo):

:O @Whitemonsterbunny17 You are actually in math ;3

whitemonsterbunny17 (whitemonsterbunny17):

Yep, English was dead. xD

OpenStudy (anonymous):

sooo do any of yall know how to do this or help me with it

sammixboo (sammixboo):

I am trying to remember how to do it -o-

OpenStudy (anonymous):

ok kcool thanx

sammixboo (sammixboo):

Noooo problem B)

whitemonsterbunny17 (whitemonsterbunny17):

Yeah, same.

OpenStudy (anonymous):

ok kcool XDXD

sammixboo (sammixboo):

I think @ganeshie8 maybe could help :/ Idk \(\LARGE \bf\rlap{\color{olive}{Sammi}}{\Huge\color{black}{\; \heartsuit ~\heartsuit ~\heartsuit }}\)

OpenStudy (anonymous):

ok kcool lol

OpenStudy (aum):

Let x be the score in the third quiz. Then, \( \Large 85 \le \frac {83 + 91 + x}{3} \le 90 \)

OpenStudy (anonymous):

ok ythanx

OpenStudy (anonymous):

i got the answer noww

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

XD

whitemonsterbunny17 (whitemonsterbunny17):

Darn, I was drawing it out. @aum beat me to it. :p

OpenStudy (anonymous):

fr well thank you also

OpenStudy (aum):

:) Sorry.